a string of pearls such that 1/3 is lost and of that 1/4th is missing, remaining is 20 then actual number of pearls?

Grade:

1 Answers

Deepak Kumar Shringi
5 years ago
total x
lost x/3
missing = x/4
remaining = x-x/3-x/4= 5x/12 = 20
x=48